OpenStudy (anonymous):

f(x,y)=e^x.sin(y) how to get the value of x, if the only information we have is f(x,y)=0 ?

OpenStudy (amistre64):

well, we should know that whenever you multiply by 0 you get zero right?

OpenStudy (australopithecus):

well if f(x,y) = 0 then set the equation to zero 0 = e^(x.sin(y))

OpenStudy (australopithecus):

and solve for x and y

OpenStudy (amistre64):

e^n is never zero

OpenStudy (anonymous):

yeah, but e^x cant be zero

OpenStudy (amistre64):

is siny in te exponent?

OpenStudy (anonymous):

nope sin(y) is not a component, sorry making u guys confuse

OpenStudy (australopithecus):

sin(0) = 0 thus we can deduce that y = 0 so we have 0 = e^(x)sin(0)

OpenStudy (amistre64):

then x is any value; as long as sin(y) = 0

OpenStudy (australopithecus):

x can be any real number :\

OpenStudy (australopithecus):

stole my thunder amistre64

OpenStudy (amistre64):

*kapow* lol

OpenStudy (australopithecus):

although to be fair I knew the answer about a minute after thinking about it I was just explaining it

OpenStudy (australopithecus):

the answer is that x is any value or the value of x is indeterminable based on the information provided

OpenStudy (amistre64):

oh yeah? i knew it three days ago and was just waiting for the opportunity to give it ;)

OpenStudy (australopithecus):

damn I can't compete with that

OpenStudy (anonymous):

Yeah i was asked to find the value of x in exam. I wonder was that a mistake ?

OpenStudy (amistre64):

x = any real number

OpenStudy (australopithecus):

x = \[\mathbb{R}\]

OpenStudy (australopithecus):

you could also say a value cannot be specified based on the information provided

OpenStudy (australopithecus):

both are right answers

OpenStudy (anonymous):

Ok thanks guys =D